DQuestion 12 5 pts Everest is more than 8 000 m high. How fast would an object be moving if it could free fall to sea level after being released from an 8000-m elevation? (Ignore air resistance.) Mt. O 4 320 km/h O 43 200 km/h O 432 km/h O 1426 km/h 1 Previous Next
Answer #1

the potential energy gets converted into Kinetic energy

mgh = 0.5mv^2

v = sqrt(2gh)

= sqrt(2*9.8*8000)

= 396 m/s

1 m/s = 18/5 kh/m

396 m/s = 396*3.6 = 1426 km/h

so last option is the correct answer